4. A sample of water was alkaline to both phenolphthalein and methyl orange. 100 mL of this
water sample required 30 mL of N/50 H2SO4 for phenolphthalein end point and another 20
mL for complete neutralisation. Determine the type and extent of alkalinity present.
1
Expert's answer
2021-10-08T02:07:42-0400

H2SO4 is dibasic acid so N\50 = 0.02N

Molarity of acid = 0.02\2 = 0.01M

So mol of acid = 10mL × 0.01M

= 0.1mmol

So amount of base = 2×0.1 mmol = 0.2mmol

Conc of base = (0.2×10-3 mol × 1000)\100

=2×10-3 M

POH = -log[OH-] = -log (2×10-3) =2.7

PH + POH = 14

PH = 14 -POH = 14 -2.7 = 11.3